OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6

OpenStudy (anonymous):

6-2 = 4 4(3) = 12?

OpenStudy (asnaseer):

not quite, you are given g(x)=6, and you are also given g(x)=3(x-2) so you just need to equate the two to get: 6 = 3(x-2) and solve from there.

OpenStudy (asnaseer):

you need to find x.

OpenStudy (anonymous):

How do I find x?

OpenStudy (asnaseer):

ok, you start with: 6 = 3(x-2) divide both sides by 3 to get: 2 = (x-2) = x-2 add 2 to both sides to get: 2+2=x-2+2 4=x so x=4 is the answer you can check this by substituting x=4 into the equation for g(x)=3(x-2): g(4)=3(4-2)=3*2=6 which matches what you were given at the start.

OpenStudy (ash2326):

g(x)=3(x-2) g(x)=6 for a particular value of x, we've to find that 6=3(x-2) 6=3x-6 12=3x x=4 solution

OpenStudy (anonymous):

Ohhh I gotcha. I was only supposed to solve for x not the actual equation, so that was my mistake lol

OpenStudy (asnaseer):

yes - at least you are learning - thats the main thing. :=)